# Free Majorana Fermion Meets Gauged Ising Conformal Field Theory on the Fuzzy Sphere

### v2

@@ 7 Discussion
    Just like the density operators, the angular components of the local fermion $\eta^{(1/2)}_{lm}=\int\mathrm{d}^{2}\mathbf{r}\,\bar{Y}^{(1/2)}_{lm}(\mathbf{r})\allowbreak f^{\dagger}(\mathbf{r})b(\mathbf{r})$ could have an algebraic structure similar to that of the density operator in fermion bilinear. In particular, it would be interesting to determine whether this algebra is a deformation of the fermionic harmonic oscillator, analogous to the free scalar on the fuzzy sphere [46]. Following this line, the ground state could be described by a wavefunction ansatz resembling a coherent state of $\eta$, as in the case of the free scalar.
Understanding these features may guide us to generalise the fuzzy sphere approach to other interacting Lagrangians. The simplest type of interaction to target could be the Yukawa coupling. Another route is to engineer the spin-$1$ gauge field and then construct the critical gauge theories.
@@
Moving on to the Ising transition, we explicitly demonstrate the physics of a topological defect of a CFT on the fuzzy sphere. Moreover, the gauging of the $\mathbb{Z}_{2}$ symmetry offers opportunities to study conformal line defects in the Ising CFT. For example, the monodromy defect in 3d Ising CFT [90] is attached to a half-infinite topological plane across which a $\mathbb{Z}_{2}$ twist is inserted. After gauging the $\mathbb{Z}_{2}$ symmetry, it becomes the vortex line of the $\mathbb{Z}_{2}$ gauge field which is no longer attached to the topological plane. This means that it can be realised locally at antipodal points on the fuzzy sphere. The vortex line at the Ising transition is closely related to the $\sigma$ anyon in the proximate bPf phase: once the bulk is gapped, the vortex line becomes the $\sigma$ worldline; under radial quantisation, this appears as two $\sigma$ anyons localised at the poles. The attempt to trap and separate anyons [91] in the $\nu=5/2$ FQH state may hint us how to realise the vortex line in the Ising CFT trap anyons in the bPf phase. Another interesting related question to explore is the self-dual confinement transition in the $\mathbb{Z}_{2}$ topological order [92]: While condensing $e$ or $m$ alone in the $\mathbb{Z}_{2}$ TO drives an Ising type transition to the trivially gapped phase, the transition condensing the $e$ and $m$ anyon at the same time in a self-dual manner may fall in a different universality. Its nature attracts wide attention, and a few critical exponents have been computed in lattice models through Monte Carlo [93]. This question could be potentially addressed on the fuzzy sphere, as the $\mathbb{Z}_{2}$ topological order could be produced by coupling the $\mathrm{SU}(2)_{2}$ (i. e. bPf) to its opposite-chirality partner $\mathrm{SU}(2)_{-2}$ and condensing certain types of anyons.
Critical theories with a discrete symmetry gauged also exist in other quantum Hall systems. One example is the direct continuous transition between $\nu_{f}=1$ fIQH state and $\nu_{b}=1/4$ bosonic Laughlin state, which was previously studied theoretically [94] and numerically [95]. These two phases can respectively be identified as the $\nu_{K}=2$ TSC and the $\mathbb{Z}_{2}$-gauged $\nu_{K}=2$ TSC in the Kitaev’s 16-fold way language. The $1/4$ bosonic Laughlin state, described by a $\mathrm{U}(1)_{-4}$ Chern-Simons theory, has four types of anyons with statistical angles $\theta_{k}=\pi k^{2}/4$ ($k=0,1,2,3$). The condensation of the fermionic $k=2$ anyon combining with the local fermion confines the other anyons and drive the transition towards fIQH, a scenario analogous to the TSC-bPf transition in our setup. More generally, we conjecture that the $\nu_{f}=1$ fIQH to $\nu=1/k^{2}$ Laughlin state could be described by a gauged Potts universality at $k=3$ or a $\mathrm{O}(2)$ Wilson-Fisher coupled to a $\mathbb{Z}_{k}$ gauge field at higher $k$.
@@ Acknowledgements.
We would like to thank Jaume Gomis, Zixiang Hu, Ryan Lanzetta, Zlatko Papić, Neville Joshua Rajappa, Subir Sachdev, Chong Wang, Zechuan Zheng, Wei Zhu, and Yijian Zou for fruitful discussions. Z. Z. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. The numerical calculations are done using the package FuzzifiED [72].
@@
Upon preparing this manuscript, we become aware of a parallel study [96] that investigates the gauged Majorana CFT (which is bosonic) in a different setup of quantum Hall bilayer transition.
@@ Appendix A Edge CFT Character and Entanglement Spectrum Degeneracy Count
$f$-wave chiral topological superconductor (TSC)
@@
$\chi^{(\text{TSC})}(q,y)$ $=\prod_{Q=0,\pm 1}q^{1/48}\prod_{k\in\mathbb{Z}+1/2}(1+y^{Q}q^{k})$
@@
is described by $\mathrm{SU}(2)_{2}$. As $\mathrm{SU}(2)$ is a double covering of $\mathrm{SO}(3)$, $\mathrm{SO}(3)_{1}\cong\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}$, i. e., the $\mathrm{SU}(2)_{2}$ can be regarded as gauging the $\mathbb{Z}_{2}$ fermion parity symmetry from $\mathrm{SO}(3)_{1}$. Correspondingly, on its edge
$\chi^{(\text{bPf})}(q,y)$ $=\tfrac{1}{2}\left[\chi^{(\text{TSC})}(q,y)\pm(q^{1/2}\to-q^{1/2})\right]$
@@ Appendix B Lagrangian Description of the Gauged Ising Transition
In this section, we construct a Lagrangian that captures the transition between the bPf phase described by $\mathrm{SU}(2)_{2}$ and the TSC phase described by $\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}\cong\mathrm{SO}(3)_{1}$. We consider a $\mathrm{U}(2)$ gauge field $\alpha+\tfrac{1}{2}a\,\mathbb{I}_{2}$ where $\alpha$ is the traceless part and $a$ is the trace part.
$S=\int\frac{2}{4\pi}\operatorname{tr}\left(\alpha\,\mathrm{d}\alpha+\tfrac{2}{3}\alpha^{3}\right)+\frac{1}{4\pi}a\,\mathrm{d}a+\frac{2}{2\pi}b\,\mathrm{d}a+\frac{1}{2\pi}A\,\mathrm{d}a+\frac{1}{4\pi}A\,\mathrm{d}A\\ +\mathrm{d}^{3}x\left[\left((\partial_{\mu}-ia_{\mu})\phi\right)^{2}+\frac{m^{2}}{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}\right]$ (17)
@@
    When $m^{2}<m_{c}^{2}$, the scalar field is condensed, this confines the $\mathbb{Z}_{2}$ gauge field. The resulting theory is $\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}=\mathrm{SO}(3)_{1}$, which describes the $f$-wave TSC.
@@ Appendix C Realisation of Local Spinor Operators on the Fuzzy Sphere
$\begin{pmatrix}\chi^{-1/2}(0)\\ 0\end{pmatrix}|\mathbb{I}\rangle=|\chi_{l=1/2,m=-1/2}\rangle,\qquad\begin{pmatrix}0\\ \chi^{1/2}(0)\end{pmatrix}|\mathbb{I}\rangle=|\chi_{l=1/2,m=1/2}\rangle.$
@@
$\mathbf{e}_{(-1/2)}=\begin{pmatrix}u\\ -v\end{pmatrix},\qquad\mathbf{e}_{(+1/2)}=\begin{pmatrix}\bar{v}\\ \bar{u}\end{pmatrix}$. (21)
@@
$\bar{\mathbf{e}}^{(s)}\mathbf{e}_{(s^{\prime})}=\delta^{s}_{s^{\prime}},\qquad\bar{\mathbf{e}}^{(s)}=(-1)^{-s}\mathbf{e}_{(-s)}^{\mathrm{T}}\mathbf{g},\quad g_{mm^{\prime}}=\delta_{m+m^{\prime},0}(-1)^{-m}.$ (22)
@@
$\langle\chi^{a}(x_{1})\chi^{b}(x_{2})\rangle=\frac{\sigma_{\mu}^{ab}x_{12}^{\mu}}{(x_{12}^{2})^{3/2}},$ (25)
where $x_{1,2}\in\mathbb{R}^{3}$, $x_{12}=x_{1}-x_{2}$, $a,b$ are spinor indices, and $\sigma$ are the Pauli matrices. It can be equivalently written as
@@
$=(-1)^{l}D^{l}_{-s,s}(\psi^{\prime},\pi-\gamma_{12},\phi^{\prime})$
$=(-1)^{l}e^{is(\phi^{\prime}-\psi^{\prime})}d_{-s,s}^{l}(\pi-\gamma_{12})$
$=e^{is(\phi^{\prime}-\psi^{\prime})}d_{-s,s}^{l}(\gamma_{12})$
$=e^{is(\phi^{\prime}-\psi^{\prime})}\frac{2l+1}{4\pi}\left(\sin\frac{\gamma_{12}}{2}\right)^{2s}P_{l-s}^{(2s,0)}(\cos\gamma_{12}).$ (32)
In the last line, we only calculate for positive $s$, and negative $s$ follows similarly. There is an overall phase that depends on the relative position of the two points, so we focus on the amplitude. In the third last line, we have used the superposition of rotations, and $\gamma_{12}$ is the angular distance between the two points. The $P_{n}^{(\alpha,0)}$ are the Jacobi polynomials. They are defined on the inteval [1,1][-1,1] and has orthonormal relation
@@
$C_{\chi}(\gamma_{12})$ $=\langle 0|\chi_{(1/2)}(\mathbf{r}_{1})\chi^{(1/2)}(\mathbf{r}_{2})|0\rangle=\sum_{l}C_{\chi,l}\frac{2l+1}{4\pi}\sin\frac{\gamma_{12}}{2}P_{l-1/2}^{(1,0)}(\cos\gamma_{12})$ (34)
$C_{\chi,l}$ $=2\pi\int\sin\gamma_{12}\mathrm{d}\gamma_{12}\,C_{\chi}(\gamma_{12})\,\sin\frac{\gamma_{12}}{2}P_{l-1/2}^{(1,0)}(\cos\gamma_{12}).$ (35)
For $C_{\chi}(\gamma_{12})=(2\sin\gamma_{12}/2)^{-2}$, the decomposed components are
@@
$\eta^{(1/2)}(\mathbf{r})=f^{\dagger}(\mathbf{r})b(\mathbf{r}),\qquad\eta^{(-1/2)}(\mathbf{r})=f^{\dagger}(\mathbf{r})b(\mathbf{r})$ (36)

### v3

@@ 9 Discussion
    Analogous to the density operators, the angular components of the local fermion $\eta^{(1/2)}_{lm}=\int\mathrm{d}^{2}\mathbf{r}\,\bar{Y}^{(1/2)}_{lm}(\mathbf{r})\allowbreak f^{\dagger}(\mathbf{r})b(\mathbf{r})$ may exhibit an algebraic structure similar to that of the density operators. It would be particularly interesting to determine whether this algebra is a deformation of the fermionic harmonic oscillator, analogous to the free scalar on the fuzzy sphere [35]. Following this line, the ground state could admit a description by a wave-function ansatz resembling a coherent state of $\eta$, as in the case of the free scalar.
The small orbital-space entanglement entropy suggests that DMRG calculations could access larger system sizes and provide more definitive answers to these questions. Understanding these features may guide us to generalise the fuzzy-sphere approach to other interacting Lagrangians, e. g. the gauge theories that require the engineering of spin-$1$ gauge fields.
Realising renormalisable Lagrangians within the fuzzy-sphere framework will allow for a systematic study of three-dimensional quantum field theories in the deep IR. This approach may address questions such as the IR fate of a given Lagrangian, and the duality between different field-theory Lagrangians [95, 96, 97]. For example, the free Majorana fermion that we have studied is conjectured to admit a dual Lagrangian description as the $\mathrm{SO}(N)_{1}$ Chern-Simons gauge theory coupled to a critical scalar in the $\mathrm{SO}(N)$ vector representation [81, 82, 83]. A fuzzy sphere realisation of the latter would allow numerical tests of this duality which go beyond comparisons of the expected phase diagrams, such as those associated with fIQH-MQH transitions.
The correspondence with Lagrangians also provides an intuitive way to design models and understand the physical meaning of the tuning parameters on the fuzzy sphere. By assigning each Majorana fermion and each real scalar an ‘excitation’ flavour in addition to a ‘reference’ flavour, various field theories can be constructed in a similar way to the super-Ising theory, including
  • The $\mathrm{O}(N_{f})$ free Majorana fermion and Gross-Neveu-Yukawa theory [90, 98],
  • the $\mathscr{N}=2$ Wess-Zumino model [99] — or equivalently the GNY theory of a Dirac fermion coupled to a complex scalar, and
  • the generalisation of Ginzburg-Landau description of a class of non-unitary minimal models to three space-time dimensions [100].
Nevertheless, our current realisation of the super-Ising theory still requires the fine tuning of three relevant operators due to the absence of the parity symmetry or manifest supersymmetry. In the fuzzy-sphere models, parity is usually implemented as the particle-hole symmetry combined with certain flavour transformations. This is not directly applicable for models with microscopic bosons. Finding a way to implement the parity for the boson-fermion-mixed models would greatly simplify targeting the fermionic CFTs. Besides that, several further questions regarding the realisation of super-Ising theory deserve attention, e. g., how the correspondence (22) between the CFT and fuzzy-sphere operators holds numerically, whether it is possible to construct the emergent super-conformal generators $Q^{\alpha}$ and $S^{\alpha}$ in analogy with the conformal generators $P^{\mu}$ and $K^{\mu}$ [30, 31], how the reference-state overlap and the small orbital-space entanglement hold for the super-Ising SCFT, etc.
Instead of letting the supersymmetry emerge, one can also try to implement part of the fermionic charges of the super-conformal symmetry at the microscopic level. The idea of fuzzy super-sphere has indeed been explored over decades [101, 102, 103]. The super-sphere $S^{2|2}$ is embedded in a super-space with three regular and two Grassmannian co-ordinates and is equipped with the $\mathrm{OSp}(1|2)$ super-rotation symmetry [104, 105]. We discuss this construction in more detail in Appendix D. The LLL on the super-sphere with a $4\pi q$ super-monopole carries a super-spin-$q$ representation, which branches into two $\mathrm{SU}(2)$ multiplets with spin-$q$ and $(q-\tfrac{1}{2})$ of opposite Grassmannian parity, in direct correspondence with our set-up. The local density-density interaction on the LLL also takes a similar form as the Hamiltonian (3) for the free Majorana fermion. However, the embedding of $\mathrm{SU}(2)$ rotations in $\mathrm{Sp}(4,\mathbb{R})$ conformal transformations does not lift to an embedding of the super-rotation group $\mathrm{OSp}(1|2)$ into the $\mathscr{N}=1$ super-conformal group $\mathrm{OSp}(1|4)$. An embedding $\mathrm{OSp}(1|2)\subset U(1|2)\subset\mathrm{OSp}(2|4,\mathbb{R})$ into the $\mathscr{N}=2$ superconformal group does exist, though. A more serious problem is that the super-rotation group $\mathrm{OSp}(1|2)$ does not admit any finite-dimensional non-trivial representation with a positive-definite inner product. If one is willing to relax the requirement of unitarity and allow for an indefinite inner product, this embedding may provide a route to engineer $\mathscr{N}=2$ SCFTs: as the rotation group is enhanced to a conformal group in the IR, the $\mathrm{OSp}(1|2)$ fermionic super-rotation generators that mix the bosonic and fermionic orbitals must also become part of a larger symmetry super-group and $\mathrm{OSp}(2|4,\mathbb{R})$ is a natural candidate.
Besides supersymmetric theories, Chern-Simons-matter theories with fermionic content is also an interesting target. For example, the confinement transition of $\nu=1/3$ Laughlin state which describes its transition with a trivially gapped state, has $2\pi$-monopole with spin-$3/2$ [106] which may be enhanced to the super-current in a $\mathscr{N}=2$ supersymmetry, but the dual supersymmetric theory is expected to be gapped [107]. Another series of examples are the QED with Chern-Simons coupling [22], which describe transitions between Jain states [108]. For example, the $N_{f}=3$ QED describes the transition between the states with fillings $\nu=1/3$ and $\nu=2/3$, and has $4\pi$-monopoles with spin-$1/2$ and $3/2$. The fuzzy sphere is a promising platform to study these theories, as the realisation of the proximate topological orders are natural with the fractional quantum Hall set-up, e. g., the Laughlin states [109] with filling $\nu=1/k$ realise the Abelian topological orders captured by Chern-Simons theories $\mathrm{U}(1)_{-k}$, the Jain sequence [108] with $\nu=p/(mp+1)$ realises the Abelian topological orders with composite fermion descriptions, and the Read-Rezayi sequence with $\nu=p/(mp+2)$ [110], as a natural extension of the Moore-Read states [58], realises in general a series of non-Abelian topological orders. Another noticeable direction is to generalise the gauged Ising transition as $\mathrm{SU}(2)_{1}$ coupled to an adjoint fermion to other adjoint QCDs [84], which may describe more exotic phase transitions.
@@
Moving on to the Ising transition, we explicitly demonstrate the physics of a topological defect of a CFT on the fuzzy sphere. Moreover, the gauging of the $\mathbb{Z}_{2}$ symmetry offers opportunities to study conformal line defects in the Ising CFT. For example, the monodromy defect in 3d Ising CFT [111] is attached to a half-infinite topological plane across which a $\mathbb{Z}_{2}$ twist is inserted. After gauging the $\mathbb{Z}_{2}$ symmetry, it becomes the vortex line of the $\mathbb{Z}_{2}$ gauge field which is no longer attached to the topological plane. This means that it can be realised locally at antipodal points on the fuzzy sphere. The vortex line at the Ising transition is closely related to the $\sigma$ anyon in the proximate bPf phase: once the bulk is gapped, the vortex line becomes the $\sigma$ worldline; under radial quantisation, this appears as two $\sigma$ anyons localised at the poles. The attempt to trap and separate anyons [112] in the $\nu=5/2$ FQH state may give an hint on how to realise the vortex line in the Ising CFT trap anyons in the bPf phase. Another interesting related question to explore is the self-dual confinement transition in the $\mathbb{Z}_{2}$ topological order [113]: While condensing $e$ or $m$ alone in the $\mathbb{Z}_{2}$ TO drives an Ising type transition to the trivially gapped phase, the transition condensing the $e$ and $m$ anyon at the same time in a self-dual manner may fall in a different universality. Its nature attracts wide attention, and a few critical exponents have been computed in lattice models through Monte Carlo [114]. This question could be potentially addressed on the fuzzy sphere, as the $\mathbb{Z}_{2}$ topological order could be produced by coupling the $\mathrm{SU}(2)_{2}$ (i. e. bPf) to its opposite-chirality partner $\mathrm{SU}(2)_{-2}$ and condensing certain types of anyons.
@@ Acknowledgements.
We would like to thank Jaume Gomis, Zixiang Hu, Ryan Lanzetta, Zlatko Papić, Neville Joshua Rajappa, Subir Sachdev, Chong Wang, Ying-Hai Wu, Zechuan Zheng, Wei Zhu, and Yijian Zou for fruitful discussions. Z. Z. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) through Discovery Grants. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Industry Canada and by the Province of Ontario through the Ministry of Colleges and Universities. The numerical calculations are done using the package FuzzifiED [76].
@@
Upon preparing this manuscript, we became aware of a parallel study [115] that investigates the gauged Majorana CFT (or Majorana, which is bosonic) in a different set-up of quantum Hall bilayer transition. The Section 8 on the super-Ising SCFT was added during a revision on January 2026. It appears after the first realisation of supersymmetry on the fuzzy sphere [116] in the gauged super-Ising universality (or super-Ising, which is bosonic).
@@ Appendix A Edge CFT Character and Entanglement Spectrum Degeneracy Count
Majorana quantum Hall (MQH) with $\nu_{K}=3$
@@
$\chi^{(\text{MQH})}(q,y)$ $=\prod_{Q=0,\pm 1}q^{1/48}\prod_{k\in\mathbb{Z}+1/2}(1+y^{Q}q^{k})$
@@
is described by $\mathrm{SU}(2)_{2}$. As $\mathrm{SU}(2)$ is a double covering of $\mathrm{SO}(3)$, $\mathrm{SO}(3)_{1}\cong\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}$, i. e., the $\mathrm{SU}(2)_{2}$ can be regarded as gauging the $\mathbb{Z}_{2}$ fermion-parity symmetry from $\mathrm{SO}(3)_{1}$. Correspondingly, on its edge
$\chi^{(\text{bPf})}(q,y)$ $=\tfrac{1}{2}\left[\chi^{(\text{MQH})}(q,y)\pm(q^{1/2}\to-q^{1/2})\right]$
@@ Appendix B Lagrangian Description of the Gauged Ising Transition
In this section, we construct a Lagrangian that captures the transition between the bPf phase described by $\mathrm{SU}(2)_{2}$ and the MQH phase described by $\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}\cong\mathrm{SO}(3)_{1}$. We consider a $\mathrm{U}(2)$ gauge field $\alpha+\tfrac{1}{2}a\,\mathbb{I}_{2}$ where $\alpha$ is the traceless part and $a$ is the trace part.
$S=\int\frac{2}{4\pi}\operatorname{tr}\left(\alpha\,\mathrm{d}\alpha+\tfrac{2}{3}\alpha^{3}\right)+\frac{1}{4\pi}a\,\mathrm{d}a+\frac{2}{2\pi}b\,\mathrm{d}a+\frac{1}{2\pi}A\,\mathrm{d}a+\frac{1}{4\pi}A\,\mathrm{d}A\\ +\mathrm{d}^{3}x\left[\left((\partial_{\mu}-ia_{\mu})\phi\right)^{2}+\frac{m^{2}}{2}\phi^{2}-\frac{\lambda}{4!}\phi^{4}\right]$ (28)
@@
    When $m^{2}<m_{c}^{2}$, the scalar field is condensed, this confines the $\mathbb{Z}_{2}$ gauge field. The resulting theory is $\mathrm{SU}(2)_{2}/\mathbb{Z}_{2}=\mathrm{SO}(3)_{1}$, which describes the $\nu_{K}=3$ MQH.
@@ Appendix C Realisation of Local Spinor Operators on the Fuzzy Sphere
$\begin{pmatrix}\chi^{-1/2}(0)\\ 0\end{pmatrix}|\mathbb{I}\rangle$ $=|\chi_{l=1/2,m=-1/2}\rangle,$ $\begin{pmatrix}0\\ \chi^{1/2}(0)\end{pmatrix}|\mathbb{I}\rangle=|\chi_{l=1/2,m=1/2}\rangle.$
@@
$\mathbf{e}_{(-1/2)}$ $=\begin{pmatrix}u\\ -v\end{pmatrix}$, $\mathbf{e}_{(+1/2)}$ $=\begin{pmatrix}\bar{v}\\ \bar{u}\end{pmatrix}$. (32)
@@
$\bar{\mathbf{e}}^{(s)}\mathbf{e}_{(s^{\prime})}$ $=\delta^{s}_{s^{\prime}},$ $\bar{\mathbf{e}}^{(s)}$ $=(-1)^{-s}\mathbf{e}_{(-s)}^{\mathrm{T}}\mathbf{g},$ $g_{mm^{\prime}}$ $=\delta_{m+m^{\prime},0}(-1)^{-m}.$ (33)
@@
$\langle\chi^{\alpha}(x_{1})\chi^{\beta}(x_{2})\rangle=\frac{\gamma_{\mu}^{\alpha\beta}x_{12}^{\mu}}{(x_{12}^{2})^{3/2}},$ (36)
where $x_{1,2}\in\mathbb{R}^{3}$, $x_{12}=x_{1}-x_{2}$, $a,b$ are spinor indices, and $\gamma$ are the Pauli matrices. It can be equivalently written as
@@
$=(-1)^{l}D^{l}_{-s,s}(\psi^{\prime},\pi-\theta_{12},\phi^{\prime})$
$=(-1)^{l}e^{is(\phi^{\prime}-\psi^{\prime})}d_{-s,s}^{l}(\pi-\theta_{12})$
$=e^{is(\phi^{\prime}-\psi^{\prime})}d_{-s,s}^{l}(\theta_{12})$
$=e^{is(\phi^{\prime}-\psi^{\prime})}\frac{2l+1}{4\pi}\left(\sin\frac{\theta_{12}}{2}\right)^{2s}P_{l-s}^{(2s,0)}(\cos\theta_{12}).$ (43)
In the last line, we only calculate for positive $s$, and negative $s$ follows similarly. There is an overall phase that depends on the relative position of the two points, so we focus on the amplitude. In the third last line, we have used the superposition of rotations, and $\theta_{12}$ is the angular distance between the two points. The $P_{n}^{(\alpha,0)}$ are the Jacobi polynomials. They are defined on the inteval [1,1][-1,1] and has orthonormal relation
@@
$C_{\chi}(\theta_{12})$ $=\langle 0|\chi_{(1/2)}(\mathbf{r}_{1})\chi^{(1/2)}(\mathbf{r}_{2})|0\rangle=\sum_{l}C_{\chi,l}\frac{2l+1}{4\pi}\sin\frac{\theta_{12}}{2}P_{l-1/2}^{(1,0)}(\cos\theta_{12})$ (45)
$C_{\chi,l}$ $=2\pi\int\sin\theta_{12}\mathrm{d}\theta_{12}\,C_{\chi}(\theta_{12})\,\sin\frac{\theta_{12}}{2}P_{l-1/2}^{(1,0)}(\cos\theta_{12}).$ (46)
For $C_{\chi}(\theta_{12})=(2\sin\theta_{12}/2)^{-2}$, the decomposed components are
@@
$\eta^{(1/2)}(\mathbf{r})$ $=f^{\dagger}(\mathbf{r})b(\mathbf{r}),$ $\eta^{(-1/2)}(\mathbf{r})$ $=f^{\dagger}(\mathbf{r})b(\mathbf{r})$ (47)